# RaffeSDG: Random Frequency Filtering enabled Single-source Domain Generalization for Medical Image Segmentation

### v2

@@ Abstract
Deep learning models often encounter challenges in making accurate inferences when there are domain shifts between the source and target data. This issue is particularly pronounced in clinical settings due to the scarcity of annotated data resulting from the professional and private nature of medical data. Despite the existence of decent solutions, many of them are hindered in clinical settings due to limitations in data collection and computational complexity. To tackle domain shifts in data-scarce medical scenarios, we propose a Random frequency filtering enabled Single-source Domain Generalization algorithm (RaffeSDG), which promises robust out-of-domain inference with segmentation models trained on a single-source domain. A frequency filter-based data augmentation strategy is first proposed to promote domain variability within a single-source domain by introducing variations in frequency space and blending homologous samples. Then Gaussian filter-based structural saliency is also leveraged to learn robust representations across augmented samples, further facilitating the training of generalizable segmentation models. To validate the effectiveness of RaffeSDG, we conducted extensive experiments involving out-of-domain inference on segmentation tasks for three human tissues imaged by four diverse modalities. Through thorough investigations and comparisons, compelling evidence was observed in these experiments, demonstrating the potential and generalizability of RaffeSDG. The code is available at https://github.com/liamheng/Non-IID_Medical_Image_Segmentation.
Index Terms:
@@
Single-source domain generalization, medical image segmentation, frequency filtering, data augmentation.
I Introduction
Deep learning-based algorithms have been widely acknowledged in numerous recent studies. However, transitioning these algorithms from controlled laboratory settings to real-world scenarios poses issues. In laboratory settings, datasets can be closed, with training (source) and test (target) data being independent and identically distributed (IID). In contrast, real-world scenarios often involve open datasets where the test data is unseen and probably non-IID with the training data [1]. This variability between training and test data is known as domain shifts, which detrimentally impact model performance [2]. Furthermore, deep learning requires an ample amount of labeled training data [3], which involves the labor-intensive and expensive task of pixel-accurate annotations in segmentation tasks. In the case of medical images, variations in imaging modalities, acquisition protocols, or patient demographics result in diverse domain shifts, which not only contribute to a scarcity of annotated data for medical image segmentation but also exacerbate the impact on clinical inference.
@@ I Introduction
Figure 1: Illustration of single-source domain generalization. (c) RaffeSDG enables generalization within (a) a single-source domain, allowing for inference over (b) a range of unseen target domains.
To mitigate the impact of domain shifts, numerous studies have explored domain adaptation (DA) and domain generalization (DG) techniques [1]. DA focuses on learning domain-invariant representations by aligning source and target domains in a feature space. However, DA relies on the assumption that both source and target data are accessible simultaneously, which is not always feasible in practice. To overcome this limitation, DG has been introduced, enabling the learning of robust models that can generalize well to any target domain, even in the absence of target data. Nevertheless, implementing DG often necessitates the availability of multiple distinct source domains, which still introduces challenges in terms of data collection. Furthermore, regardless of whether it involves simultaneously accessing source and target data or leveraging multiple source domains, concerns about privacy preservation arise, particularly in sensitive medical scenarios. As an alternative to DA and DG, single-source domain generalization (SDG) [2] has been proposed, which focuses on a worst-case scenario in DG involving out-of-domain inference from a single-source domain (shown in Fig. 1). By utilizing only a single-source domain, SDG circumvents data collection and transmission across different centers, mitigating data dependency and privacy concerns. Various SDG paradigms [4, 5] have been developed to generalize segmentation models from a single-source medical dataset across a range of unseen distributions.
Despite the dedicated efforts to deploy deep learning models in real-world scenarios, challenges persist in clinical segmentation tasks. i) The limited availability of annotations in clinical deployment prompts researchers to utilize existing datasets with similar characteristics, aiming to reduce the burden of annotation. However, this situation inevitably encounters domain shift issues. ii) While popular DA and DG algorithms tackle domain shifts, their reliance on additional data dependency imposes additional burdens in data collection and raises privacy concerns. iii) The utilization of complex feature operations and auxiliary generative networks in existing SDG paradigms often hinder their efficiency and versatility, thereby impeding their successful clinical deployment.
@@
To facilitate the application of SDG, we introduce a Random frequency filtering enabled SDG algorithm (RaffeSDG), which mitigates the computational and network training costs. RaffeSDG employs random frequency filtering and homologous sample blending to conveniently introduce randomization within the single-source domain through data augmentation. Furthermore, RaffeSDG also incorporates Gaussian filter-based structural saliency to construct a generalizable segmentor by learning robust representations in segmentation. Three human tissues imaged by four modalities are leveraged in the experiment to demonstrate the efficacy of RaffeSDG in out-of-domain inference. Our main contributions are summarised as follows:
  • An efficient SDG paradigm for medical image segmentation termed RaffeSDG is developed, to alleviate the challenges in out-of-domain inference without introducing additional data dependency and privacy concerns.
@@
  • To promote diversity within a single-source domain, frequency filter-based data augmentation is accomplished using random frequency filters and sample blending to perform domain randomization.
  • Cooperating with the augmentation, a generalizable segmentor is constructed by learning robust representations across samples through Gaussian filter-based structural saliency.
@@
II Related Work
II-A Domain adaptation and domain generalization
@@
DA encompasses both supervised and unsupervised paradigms. In supervised DA, the focus is on few-shot adaptation where only a few labeled target samples are involved in training. Nevertheless, in medical scenarios, the scarcity of annotated data makes unsupervised DA more prevalent, which aims to align the distribution of the target domain with that of the source domain by minimizing the maximal mean variability or utilizing a domain classifier. To achieve adaptation across domains, DAMAN [6] employs a domain regularizing loss to align intermediate features between the source and target domains. CSCADA [7] achieves semi-supervised DA utilizing source data through anatomical structure discrimination. Nevertheless, accessing the target domain poses a strong requirement in practice, and both paradigms suffer from a strong coupling between the source and target domains, limiting their practicality and generalizability in clinical settings.
DG serves as an alternative pipeline to tackle domain shifts [1]. In contrast to DA, the objective of DG is to learn from multiple source domains without requiring access to target domains. This reduced reliance on target data makes DG more appreciated in medical scenarios. Recently, frequency operations have gained prominence in DG algorithms to generate diverse additional training data by domain randomization [8]. FACT [9] randomizes source domains by swapping the low-frequency component between multiple source domains. Moreover, FedDG [10] improved the augmentation by employing continuous frequency space interpolation, further enhancing the realism of the generated data. Apart from augmentation techniques, learning techniques such as feature regularization and self-supervision have gained significant popularity for imposing the learning of domain-invariant representations. To learn generalizable representations, co-teacher regularization and contrastive learning were respectively leveraged in FACT [9] and FedDG [10]. Unfortunately, the requirement for multiple distinct source domains in typical DG approaches still poses difficulties in terms of data collection and preserving privacy, especially in clinical settings.
II-B Single-source domain generalization
@@
An intuitive solution of SDG is to augment the diversity of the single-source domain. To learn a generalizable model from a single-source, uncertainty-guided model generalization [11, 2] was posed, which augments the source capacity in both input and label spaces. Subsequently, attention consistency on visual corruptions (ACVC) [12] imposed the robustness to superfluous correlations in a single-source domain to generalize visual recognition models. To facilitate the augmentation within a single-source, SLAug [13] was developed to provide both global and local augmentation. Inspired by causality, Global Intensity Non-linear augmentation (GIN) and Interventional Pseudocorrelation Augmentation (IPA) were devised [14] to achieve cross-domain medical image segmentation from a single-source. Furthermore, our previous study [15] has demonstrated the effectiveness of introducing domain randomization through frequency filtering, leading to the development of FreeSDG, an SDG algorithm that utilizes diverse Gaussian filters. In contrast, SDG algorithms that aim to learn invariant representations from a single-source domain have been proposed as a complementary approach to the data augmentation paradigm. Inspired by feature normalization, Semantic-Aware Normalization (SAN) and Semantic-Aware Whitening (SAW) were developed [16] to enhance the learning of domain-invariant and discriminative features.
Despite the impressive success of existing SDG algorithms, their reliance on complex feature operations or auxiliary generative networks hinders practical efficiency and versatility. To address these limitations, we have previously proposed FreeSDG [15] as a solution. However, the naive frequency operations employed in FreeSDG [15] restrict the flexibility of data augmentation (refer to Fig. 7). Therefore, RaffeSDG further incorporates random Fourier frequency filters to introduce an efficient SDG paradigm for medical image segmentation.
III Method
@@ III-A Definition and overview
In DG tasks, the training data from source domains are denoted as $\mathbb{D}_{S}=\{(x^{s}_{n},y^{s}_{n})^{N_{s}}_{n=1}\}^{S}_{s=1}$, where $x^{s}_{n}$ refers to the $n^{th}$ sample in the $s^{th}$ source domain, $y^{s}_{n}$ is the corresponding ground-truth, and $N_{s}$ and $S$ are the domain volume and the number of source domains, respectively. Based on $\mathbb{D}_{S}$, the parameters $\theta$ of a prediction function $g_{\theta}$ are learned, which maps a sample $x$ from the image space $\mathcal{X}$ to a label $y$ in the label space $\mathcal{Y}$, i.e., $g_{\theta}:x\to y$. The main goal of DG is to impose the learned function $g_{\theta}$ generalizes well to unseen target domains $\mathbb{D}_{T}=\{(x^{t}_{n},y^{t}_{n})^{N_{t}}_{n=1}\}^{T}_{t=1}$, namely minimizing the error between the prediction $\hat{y}^{t}_{n}=g_{\theta}(x^{t}_{n})$ and the ground-truth $y^{t}_{n}$.
In pioneer studies, to ensure that $g_{\theta}$ remains robust against domain shifts between $\mathbb{D}_{S}$ and $\mathbb{D}_{T}$, it was common to use multiple source domains ($S>1$) for model training. More recently, to address the bottleneck in collecting multiple source domains, SDG has emerged to achieve DG in the worst-case scenario, where only one source domain ($S=1$) is available during training.
In our previous study [15], we validated the potential of frequency variations in domain randomization and successfully implemented SDG in medical image segmentation. However, the previous SDG algorithm had limitations in terms of data augmentation diversity. Therefore, we have introduced RaffeSDG in this study, which offers flexible data augmentation by utilizing random frequency filtering and homologous sample blending within the single-source domain. Subsequently, a Gaussian filter is further incorporated with RaffeSDG to boost the training of a generalizable segmentor.
III-B Data augmentation for domain randomization
To boost the generalization and robustness of models, SDG algorithms commonly rely on data augmentation to introduce randomization within the single-source domain. As a result, previous studies have developed various data augmentation techniques, including generative networks [2, 17, 14] and visual corruption methods [18, 12]. However, the use of generative networks comes with additional computational costs and may encounter the problem of mode collapse, while visual corruption can potentially lead to the destruction of crucial structures in medical images. To overcome these limitations, we propose an effective domain randomization approach that conveniently augments data from a single-source domain using random filters and sample blending.
III-B1 Frequency filters based on Fourier transform
Domain shifts can be bridged by aligning the low-level statistics between the source and target domains. Recently studies such as Fourier-based domain adaptation [19] and generalization [8], have proven that matching the low-frequency spectrum (LFS) in image samples enables this alignment. Consequently, exchanging or integrating the LFS of one domain with another has been widely utilized to address domain shifts [9, 10].
Inspired by the Fourier-based DA and DG techniques, we hypothesize that the use of multiple source domains is not necessary for implementing domain randomization. Instead, we propose that domain randomization can be conveniently achieved by introducing random LFS variations to samples within a single-source domain. Specifically, the desired LFS variations are introduced by applying random parameters to Fourier-based high-pass filters during the frequency filtering of image samplesas, as exhibited in Fig. 2 (a).
@@ III-B1 Frequency filters based on Fourier transform
Additionally, considering filters with sharp transitions can lead to the occurrence of ringing artifacts, the high-pass filter is implemented with the Butterworth filter, which is defined as
@@
The LFS variations can be introduced by adjusting the parameters of $D_{0}$ and $n$ (the parameter range is interpreted in Sec. IV-B). Notably, as depicted in Fig. 2 (a), distinct parameters are applied to the filter across different channels, promising the diversity of filtered samples.
@@
III-B2 Homologous sample blending
Although frequency filtering affects the global image, it lacks the ability to discriminatively alter local areas of the images. To overcome this limitation, we draw inspiration from popular augmentation techniques such as CutOut [20] and CutMix [21]. Consequently, homologous filtered samples are blended to introduce additional diversity at the sub-image level, effectively enhancing data augmentation.

### v3

@@ Abstract
Deep learning models often encounter challenges in making accurate inferences when there are domain shifts between the source and target data. This issue is particularly pronounced in clinical settings due to the scarcity of annotated data resulting from the professional and private nature of medical data. Although various cross-domain strategies have been explored, including frequency-based approaches that vary appearance while preserving semantics, many remain limited by data constraints and computational cost. To tackle domain shifts in data-scarce medical scenarios, we propose a Random frequency filtering enabled Single-source Domain Generalization algorithm (RaffeSDG), which promises robust out-of-domain inference with segmentation models trained on a single-source domain. A frequency filter-based data augmentation strategy is first proposed to promote domain variability within a single-source domain by introducing variations in frequency space and blending homologous samples. Then Gaussian filter-based structural saliency is also leveraged to learn robust representations across augmented samples, further facilitating the training of generalizable segmentation models. To validate the effectiveness of RaffeSDG, we conducted extensive experiments involving out-of-domain inference on segmentation tasks for three human tissues imaged by four diverse modalities. Through thorough investigations and comparisons, compelling evidence was observed in these experiments, demonstrating the potential and generalizability of RaffeSDG. The code is available at https://github.com/liamheng/Non-IID_Medical_Image_Segmentation.
keywords:
@@
Single-source domain generalization , Medical image segmentation , Frequency filtering , Data augmentation
journal: Biomedical Signal Processing and Control
[…]
[label1]organization=Faculty of Biomedical Engineering, Shenzhen University of Advanced Technology, city=Shenzhen, state=Guangdong, country=China
[…]
[label2]organization=The Department of Computer Science and Engineering, Southern University of Science and Technology, city=Shenzhen, state=Guangdong, country=China
[…]
[label3]organization=The Department of Biostatistics, School of Global Public Health, New York University, city=New York, state=New York, country=USA
[…]
[label4]organization=Department of Ophthalmology, Peking University People’s Hospital, city=Beijing, country=China
[…]
[label5]organization=Beijing Key Laboratory of Ocular Disease and Optometry Science, Peking University People’s Hospital, city=Beijing, country=China
1 Introduction
Deep learning algorithms have garnered widespread recognition in medical research. However, their performance often deteriorates when transitioning from controlled laboratory settings to clinical scenarios due to domain shifts—discrepancies between training (source) and test (target) data distributions caused by differences in imaging modalities, acquisition protocols, or patient demographics [zhou2022domain, peng2022out]. To mitigate the impact of domain shifts, numerous studies have explored techniques such as domain adaptation (DA) and domain generalization (DG) [zhou2022domain]. DA focuses on adapting a model trained on a source domain to a target domain, but requires access to the target domain during training, which is often impractical in open clinical datasets. DG addresses this limitation by leveraging multiple source domains to build models that generalize well on unknown domains. Despite their benefits, both DA and DG raise concerns regarding data collection and privacy, especially in clinical scenarios. Consequently, single-source domain generalization (SDG) [peng2022out] has recently been introduced to enable out-of-domain inference using only a single-source domain, thereby circumventing these concerns.
@@ 1 Introduction
Figure 1: Illustration of single-source domain generalization. (c) RaffeSDG enables generalization within (a) a single-source domain, allowing for inference over (b) a range of unknown target domains.
As a worst-case scenario of DG (shown in Fig. 1 (a) (b)), SDG typically employs augmentation techniques to enhance diversity within the single domain, such as perturbing local textures [ouyang2022causality, hu2024domain, liang2024single] or transforming global styles [su2022rethinking] to simulate potential domain variations. While these methods primarily operate in the spatial domain, recent studies have increasingly explored alternative augmentation strategies that manipulate the frequency space. Frequency-based approaches have been widely adopted in cross-domain settings for their ability to induce substantial appearance variations while preserving semantic content [cheng2023frequency], making them particularly appealing for scenarios where semantic integrity must be maintained despite distributional shifts. For instance, Yang et al. [yang2020fda] proposed a frequency-driven method by replacing the amplitude spectrum of a source image with that of a target image. Extending this idea, subsequent algorithms [xu2021fourier, liu2021feddg] encouraged feature invariance across domains undergoing amplitude spectrum transformations. More recently, previous work [li2023frequency] introduced a frequency-oriented SDG paradigm, which leverages parameterized Gaussian filters to manipulate the frequency components of source images for enhancing intra-domain diversity.
Despite advancements in generalizing deep learning models for clinical applications, challenges persist in implementing SDG, particularly the frequency-enabled paradigm. i) Pioneer SDG algorithms primarily rely on image transformation [cugu2022attention] or generative networks [peng2022out, guo2025everything, ouyang2022causality] to enhance data diversity. However, the former may offer negative augmentation, while the latter can escalate the risk of mode collapse. ii) Though previous frequency-based SDG algorithm alleviated these issues, its use of straightforward Gaussian filters is insufficient to fully diversify the source domain, necessitating more efficient frequency augmentation methods. iii) Beyond enhancing data diversity within the single-source domain, reasonably leveraging the augmented data to learn and embed generalizable representations is also significant in SDG algorithms.
@@
To facilitate the implementation of SDG, we introduce a Random frequency filtering enabled SDG algorithm (RaffeSDG), which allows generalizable segmentation models from a single-source domain without the drawbacks of negative augmentation and mode collapse. RaffeSDG employs random Fourier filtering and homologous sample blending to augment data diversity within the single-source domain (Fig. 1 (c)). Furthermore, RaffeSDG leverages structure consistency and saliency constraints to impose model robustness and generalizability. Three human tissues imaged by four modalities are leveraged in the experiment to demonstrate the efficacy of RaffeSDG in out-of-domain inference. Our main contributions are summarised as follows:
  • 1.
    A frequency-enabled SDG paradigm for medical image segmentation termed RaffeSDG is developed, to alleviate the challenges in out-of-domain inference while avoiding the risk of negative augmentation and mode collapse.
  • 2.
@@
    A frequency-based augmentation strategy is developed using random Fourier filters and sample blending to attain convenient and effective data diversity within a single-source domain.
  • 3.
    Cooperating with the augmentation strategy, constraints on structure consistency and saliency are integrated to boost the robustness and generalizability during model training.
  • 4.
@@
2 Related Work
2.1 Domain adaptation and domain generalization
@@
DA encompasses both supervised and unsupervised paradigms. In supervised DA, the focus is on few-shot adaptation where only a few labeled target samples are involved in training. Nevertheless, in medical scenarios, the scarcity of annotated data makes unsupervised DA more prevalent, which aims to align the distribution of the target domain with that of the source domain by minimizing the maximal mean variability or utilizing a domain classifier. To achieve adaptation across domains, DAMAN [mukherjee2022domain] employs a domain regularizing loss to align intermediate features between the source and target domains. CSCADA [gu2022contrastive] achieves semi-supervised DA utilizing source data through anatomical structure discrimination. AIF-SFDA [li2025aif] trains adaptive filters on unlabeled target data to disentangle domain-specific information and boost source model generalizability. Nevertheless, accessing the target domain poses a strong requirement in practice, and both paradigms suffer from a strong coupling between the source and target domains, limiting their practicality and generalizability in clinical settings.
DG serves as an alternative pipeline to tackle domain shifts [zhou2022domain, li2025federated]. In contrast to DA, the objective of DG is to learn from multiple source domains without requiring access to target domains. This reduced reliance on target data makes DG more appreciated in medical scenarios. Recently, frequency operations have gained prominence in DG algorithms to generate diverse additional training data by domain randomization [niu2025exploring]. FACT [xu2021fourier] randomizes source domains by swapping the LFS between multiple source domains. Moreover, FedDG [liu2021feddg] improved the augmentation by employing continuous frequency space interpolation, further enhancing the realism of the generated data. Apart from augmentation techniques, learning techniques such as feature regularization and self-supervision have gained significant popularity for imposing the learning of domain-invariant representations. To learn generalizable representations, co-teacher regularization and contrastive learning were respectively leveraged in FACT [xu2021fourier] and FedDG [liu2021feddg]. Unfortunately, the requirement for multiple distinct source domains in typical DG approaches still poses difficulties in terms of data collection and preserving privacy, especially in clinical settings.
2.2 Single-source domain generalization
@@
A common SDG strategy is to augment the data distribution while preserving semantic integrity, typically assuming a separation between domain-relevant (e.g., style, texture) and domain-irrelevant (e.g., content) factors. To this end, commonly adopted techniques include perturbing texture information via randomized auxiliary networks [ouyang2022causality, hu2024domain, xu2025adversarial], modulating image style through histogram-based transformations [su2022rethinking] or randomized normalization methods [liang2024single] based on adaptive instance normalization (AdaIN) [huang2017arbitrary], and applying randomized frequency-domain filtering, which enables effective augmentation without access to target domain spectral information [li2023frequency]. Feature-level constraint  [liang2024single, cho2025peer] is another prevalent strategy in SDG, aiming to improve generalization by regulating the structure of learned representations. Typical practices include applying contrastive learning to emphasize inter-domain discrepancies between the source and its augmented variants [liang2024single], enforcing Lipschitz continuity to reduce sensitivity to domain-specific perturbations [arslan2024single].
Despite the impressive success of existing SDG algorithms, their reliance on complex feature operations or auxiliary generative networks hinders practical efficiency and versatility. To address these limitations, FreeSDG [li2023frequency] has previously proposed a solution. However, the naive frequency operations employed in FreeSDG [li2023frequency] restrict the flexibility of data augmentation (refer to Fig. 6). Therefore, RaffeSDG further incorporates random Fourier frequency filters to introduce an efficient SDG paradigm for medical image segmentation.
3 Method
@@ 3.1 Definition and overview
In DG tasks, the training data from source domains are denoted as $\mathbb{D}_{S}=\{(x^{s}_{n},y^{s}_{n})^{N_{s}}_{n=1}\}^{S}_{s=1}$, where $x^{s}_{n}$ refers to the $n^{th}$ sample in the $s^{th}$ source domain, $y^{s}_{n}$ is the corresponding ground-truth, and $N_{s}$ and $S$ are the domain volume and the number of source domains, respectively. Based on $\mathbb{D}_{S}$, the parameters $\theta$ of a prediction function $g_{\theta}$ are learned, which maps a sample $x$ from the image space $\mathcal{X}$ to a label $y$ in the label space $\mathcal{Y}$, i.e., $g_{\theta}:x\to y$. The main goal of DG is to impose the learned function $g_{\theta}$ generalizes well to unknown target domains $\mathbb{D}_{T}=\{(x^{t}_{n})^{N_{t}}_{n=1}\}^{T}_{t=1}$. To ensure that $g_{\theta}$ remains robust against domain shifts between $\mathbb{D}_{S}$ and $\mathbb{D}_{T}$, it was common to use multiple source domains ($S>1$) for model training. More recently, to address the bottleneck in collecting multiple source domains, SDG has emerged to achieve DG with only one source domain ($S=1$).
Inspired by the inter-domain attributes of image frequency, we have explored using Gaussian lowpass filters to develop a frequency-enabled SDG paradigm, termed FreeSDG [li2023frequency]. Despite the progress achieved by FreeSDG [li2023frequency], Gaussian filters suffer a tendency to remove color information, leading to limited diversity in augmentation. Therefore, we introduce RaffeSDG in this study, which offers flexible data augmentation through random Fourier filtering and homologous sample blending within a single-source domain. Subsequently, structure consistency and saliency are incorporated with the augmented data to train generalizable segmentation models.
3.2 Data augmentation using random frequency filtering
Data augmentation has been commonly employed in SDG algorithms to enhance the diversity within a single-source domain. However, pioneer SDG algorithms employing image transformation [tiwary2025langdaug, cugu2022attention] or generative networks [peng2022out, guo2025everything, ouyang2022causality] suffer from negative augmentation or model collapse. To overcome these limitations, we develop a frequency-based data augmentation strategy using random Fourier filters and sample blending to diversify the single-source domain.
3.2.1 Random Fourier filtering with Butterworth filers
Considering AFS contains low-level image statistics, it has been leveraged in DA and DG to bridge domain shifts. Accordingly, we further hypothesize that diversifying a domain can be effectively achieved by adjusting the AFS of image samples. To this end, we employ Fourier filtering to enhance the diversity within the single-source domain through data augmentation with AFS adjustments, as illustrated in Fig. 2 (a).
@@ 3.2.1 Random Fourier filtering with Butterworth filers
Additionally, considering filters with sharp transitions can lead to the occurrence of ringing artifacts, the high-pass filter is implemented with Butterworth filters, which is defined as
@@
The AFS variations can be introduced by adjusting the parameters of $D_{0}$ and $n$ (the parameter range is interpreted in Sec. 4.2). Notably, as depicted in Fig. 2 (a), distinct parameters are applied to the filter across different channels, promising the diversity of filtered samples.
@@
3.2.2 Homologous sample blending
Although frequency filtering affects the global image, it lacks the ability to discriminatively alter local areas of the images. To overcome this limitation, we draw inspiration from popular augmentation techniques such as CutOut [devries2017improved] and CutMix [yun2019cutmix]. Consequently, homologous filtered samples are blended to introduce additional diversity at the sub-image level, effectively enhancing data augmentation.